Randolph Company reported pretax net income from continuing operations of $869,000 and taxable income of $580,000. The book–tax difference of $289,000 was due to a $286,000 favorable temporary difference relating to depreciation, an unfavorable temporary difference of $95,000 due to an increase in the reserve for bad debts, and a $98,000 favorable permanent difference from the receipt of life insurance proceeds.

a. Compute Randolph Company’s current income tax expense.

b. Compute Randolph Company’s deferred income tax expense or benefit.

c. Compute Randolph Company’s effective tax rate. (Round your answer to 2 decimal places.)

d. Provide a reconciliation of Randolph Company’s effective tax rate with its hypothetical tax rate of 21 percent. (Amounts to be deducted should be indicated by a minus sign. Round your percentages to 2 decimal places.)
Answer #1

a) Current Income tax payable

Pre tax net income. 869000

Favourable temporary difference. (286000)

Unfavorable temporary difference. 95000

Favourable permanent difference. (98000)

Taxable income. 580000

*21%(hypothetical tax rate)

Current income tax payable. 121800

b) Deferred income tax expense or benefit

Favourable temporary difference. (286000)

Unfavorable temporary difference. 95000

Net favourable temporary difference (191000)

*21%

Deferred tax expense (liability) (40110)

Total Income tax provision =121800+40110

=161910

c) Effective tax rate =161910/869000*100

=18.63%

d) ETR reconciliation

Income tax expense at 21% (869000*21%)=182490

less: Tax benefits from permanent difference (98000*21%)=(20580)

Income tax provision =161910

Hypothetical income tax rate. 21%

Less:Tax benefits from permanent difference (20580/869000)*100. = (2.37%)

effective tax rate = 18.63%